OpenStudy (anonymous):

What value of a makes this equation true? 9/10 =a/30 3 90 27 12

OpenStudy (knightofpentacles):

9x30= 270 10/270=27

OpenStudy (knightofpentacles):

want me to demonstrate?

OpenStudy (anonymous):

The city planning committee is composed of 12 women and 14 men. What fraction of the committee is composed of women? 6/13 7/13 6/7 1/12

OpenStudy (knightofpentacles):

6/7

OpenStudy (knightofpentacles):

2/12= 6 2/14=7

OpenStudy (knightofpentacles):

answer is 6/7

OpenStudy (knightofpentacles):

?

OpenStudy (anonymous):

A pizza is cut into eight slices. The slices are all the same size. Carolina ate two slices. What fraction of the pizza did Carolina eat? 1/2 1/8 1/3 1/4

OpenStudy (knightofpentacles):

8 slices she ate 2 6/8 in simplified form 3/4

OpenStudy (anonymous):

so which one is it of one of those answer I put on my post

OpenStudy (knightofpentacles):

does the question asked you to simplified form?

OpenStudy (anonymous):

no

OpenStudy (knightofpentacles):

maybe its asking you to do it automatically

OpenStudy (knightofpentacles):

just in case

OpenStudy (knightofpentacles):

sometimes they wont mention to simplify but if even so if you were to put an answer as 6/8 sometimes it could be wrong when you don't simplify all the way down write down the process i said in that order im sure you get it right

OpenStudy (anonymous):

ok I got it

OpenStudy (anonymous):

27